# Misinterpreting Spin Precession as Orbital Eccentricity in Gravitational-Wave Signals

### v1

@@ I Introduction
This past year marked the tenth anniversary of the first detection of gravitational waves (GW) from the merger of a binary black hole (BBH). The loud chirp of GW150914 LIGOScientific:2016aoc was a landmark discovery that propelled the community into various streams of scientific investigation. Since that first discovery, the The LIGO-Virgo-Kagra (LVK) collaboration has reported 218 events in the four gravitational-wave catalogs LIGOScientific:2018mvr; LIGOScientific:2020ibl; LIGOScientific:2021usb; KAGRA:2021vkt; LIGOScientific:2025slb. As the number of observed events increases, the likelihood of encountering systems with non-generic or exotic features, such as orbital eccentricity and spin-induced precession also increases Hoy:2024wkc; Gumbel1958_EVT.
These non-generic features, in turn, offer insight into one of the key open questions in the field: the identification of BBH formation mechanisms; How and when the binary formed, and in what environment. Binary properties like eccentricity or spins tilted below the orbital plane can indicate that the binary underwent dynamical interactions at some point in its lifetime, e.g. Samsing:2017xmd; Rodriguez:2016vmx. Thus, accurate and reliable measurement of these features is of great importance. Occasionally, the GWs observed by the detectors will capture only the final cycles of a binary system’s inspiral and merger. That can give rise to the challenge of using limited information, sometimes only a handful of cycles in the sensitive band, to reconstruct as much of the binary’s history as possible.
The GW observations have revealed a potential population of massive BHs LIGOScientific:2025pvj, with examples like GW190521 having an estimated total mass of $\sim 150\,M_{\odot}$ LIGOScientific:2020iuh; LIGOScientific:2020ufj and GW231123 with $\sim 240\,M_{\odot}$ LIGOScientific:2025slb; LIGOScientific:2025rsn. These massive signals have few GW cycles in band, only $\sim 4$–5, making robust interpretation challenging Fumagalli:2024gko. A small number of cycles poses a challenge in placing constraints on the source properties, particularly parameters such as eccentricity and in-plane spin components. The initial LVK analysis of GW190521 included quasi-circular precessing templates LIGOScientific:2020iuh; LIGOScientific:2020ufj, and found mild evidence for the imprint of precession; however, it did not include any with eccentricity. Since then, multiple studies have reanalyzed the event; however, its origin remains ambiguous. Work by various groups has shown that the signal could have been emitted by an eccentric source Romero-Shaw:2020thy; Gayathri2022, a precessing source Miller:2023ncs, or more exotic compact objects such as boson stars CalderonBustillo:2020fyi or cosmic strings Cuceu:2025fzi. These competing propositions highlight the challenge of positively inferring eccentricity from short signals, and the possibility that precession and eccentricity can be misidentifed with limited models.
@@
Eccentricity and spin-precession are relatively well-studied parameters individually. Both these effects cause a modulation in amplitude and frequency in the GW signal. However, these effects generally operate on different timescales. While eccentricity produces modulations at an orbital timescale, precession can cause modulations on longer timescales, spanning many orbits. Superficially, these distinct timescales can make the waveforms produced by systems displaying eccentric or precessing dynamics appear starkly different. Nevertheless, as illustrated in Fig. 1, there exists regions of parameter space in which their modulations overlap in amplitude and phase in a way that can make them qualitatively similar and potentially degenerate. Romero-Shaw et al. Romero-Shaw:2022fbf reported that a sufficient number of cycles in-band provides enough information to distinguish between these effects; however, this is not the case for a GW190521-like signal, i.e., signals with few cycles in-band. Contrary to expectations, Xu et al. Xu2023 focused on heavier systems but reported no correlation between eccentricity and precession in the specific numerical relativity (NR) systems they studied. More recently, Divyajyoti et al. Divyajyoti:2025cwq expanded this picture by exploring parameter biases for eccentric systems with and without precessional effects when analyzed with waveform models with incomplete physics. They find that for systems exhibiting both these effects simultaneously, an eccentric aligned spin model is preferred over a quasi-circular precessing model. This highlights how waveform systematics can mimic an underlying physical degeneracy. Overall, the existence and strength of this degeneracy across the broader parameter space remain largely unclear, despite differing interpretations of events like GW190521 when using different models.
@@
Meanwhile, the astrophysical importance of measuring eccentricity has driven developments in the modeling of GWs from eccentric systems, as well as assessing the impact of mismodeling on PE. A number of state-of-the-art models exist that treat precession and eccentricity separately (see Sec. II). Such models show that neglecting eccentricity when present leads to biases in interpretations, e.g. OShea:2021faf; Favata:2021vhw; Divyajyoti:2023rht; Saini:2022igm; Zeeshan:2024ovp. They have been used to supplement quasi-circular, precessing inferences of detected BBH GW events with additional eccentric analyses, e.g. Romero-Shaw:2021ual; Iglesias:2022xfc; Gupte:2024jfe; Planas:2025jny; Romero-Shaw:2025vbc, which in some cases suggest non-negligible eccentricities for a subset of events. Finally, a small number of waveform models have been developed which incorporate both eccentricity and precession Klein:2018ybm; Gamba:2024cvy; Morras:2025xfu; Morras:2025nlp. These are either limited to only the inspiral phase, or are relatively new models which remain computationally expensive for use in full-scale PE studies.
@@
This paper is organized as follows. We describe the waveform model and associated conventions in Sec. II. In Sec. III, we present the setup and results of the preliminary mismatch study that informs our parameter estimation section. In Sec. IV, we detail the setup and results for PE for the identified injection set. Finally, in Sec. V, we discuss the implications and robustness of our results and outline directions for future work. Throughout this work we set $G=c=1$, quote detector frame rather than (redshifted) source-frame masses LIGOScientific:2025hdt, and adopt the NR convention that the mass ratio is greater than unity, $q\geq 1$.
@@ II Waveform model
A number of state-of-the-art waveform models exist with differing conventions, approximations, and physics content. This includes phenomenological models 2021PhRvD.103j4056P; 2025PhRvD.111j4019C; PhysRevD.104.124027; PhysRevD.109.063012; 2025arXiv250702604H; PhysRevD.105.084039; PhysRevD.105.084040, effective-one-body (EOB) models from the SEOB family PhysRevD.95.044028; PhysRevD.98.084028; PhysRevD.102.044055; PhysRevD.101.124040; PhysRevD.104.124087; 2023PhRvD.108l4035P and the TEOB family 2020PhRvD.102b4077N; PhysRevD.104.104045; PhysRevD.103.024014; PhysRevD.106.024020; PhysRevD.110.024031, and the NR surrogate models 2019PhRvR...1c3015V; 2019PhRvD..99f4045V. Many of these model families also have there own associated eccentric model 2025arXiv250313062D; Gamboa2024; 2022PhRvD.105d4035R; 2020PhRvD.101j1501C; 2021PhRvD.103f4022I. For this work, we use SEOBNRv5 waveforms that are based on the EOB formalism. EOB models approximate the full inspiral–merger–ringdown regime. We adopt SEOBNRv5EHM Gamboa2024 as our non-precessing eccentric model and SEOBNRv5PHM Ramos-Buades:2023ehm; 2023PhRvD.108l4035P as our quasi-circular precessing model. Using waveforms from the same model family minimizes potential model-related systematics when comparing the two. Both models support higher-order modes. The modes used for SEOBNRv5EHM are {(2,±2), (2,±1), (3,±3), (3,±2), (4,±4), (4,±3)}, and for SEOBNRv5PHM are {(2,±2), (2,±1), (3,±3), (3,±2), (4,±4), (4,±3), (5,±5)} in the co-precessing frame. SEOBNRv5EHM allows for explicit sampling in relativistic anomaly $\zeta$ at the reference frequency, an essential parameter in the relativistic definition of eccentricity.
@@ III.1 Parameter choices
Mismatches 1999PhRvD..60b2002O are conventionally used to quantify the similarity between two waveforms, for example when calibrating analytical models against post-Newtonian or NR waveforms Boyle:2007ft; Hannam:2007ik; Buonanno:2009zt; Ajith:2007qp. In this work, we use the mismatch between waveforms generated by physically distinct systems to assess the similarity between eccentric and precessing signals. This serves as a relatively inexpensive diagnostic for the expected behavior of full PE. This is a suboptimal approach, as we can only maximize over a few parameters while keeping several extrinsic and intrinsic parameters fixed.
@@
where $h_{p}$ is the precessing signal, $h_{e}$ is the eccentric signal and $\langle\cdot|\cdot\rangle$ is the standard noise-weighted inner product (Appx. B). For the mismatch we maximize the inner product over the difference in coalescence time $\Delta t$, coalescence phase $\Delta\phi$, and polarization angle $\Delta\psi$ between the signals. We fix the same total mass $M_{\mathrm{tot}}$ and mass ratio $q$ for both signals. In all cases we fix the inclination to $\pi/3$ radians, a value intermediate between the special face-on ($\iota=0$) and edge-on ($\iota=\pi/2$) cases. This inclination value also coincides with the peak of the expected observed distribution of inclinations Schutz:2011tw. Meanwhile, as the two models support different physical effects, our choices of spins and eccentricity parameters necessarily differ.
@@
The number of cycles in band is a significant factor in identifying eccentricity Romero-Shaw:2022fbf. For a given total mass, the mass ratio and number of cycles are inversely related Cutler:1994ys (see Fig. 2); therefore we expect that misidentification of eccentricity is more likely in short duration signals. As a result, we focus on high total mass systems having $M_{\text{tot}}\in[200M_{\odot},300M_{\odot}]$ (detector-frame), while varying the mass ratios $q\in[1,3]$ to control the number of cycles in band.
@@ IV.1 PE methodology
We use Bayesian inference on injected signals using the parameter inference pipeline RIFT Pankow:2015cra; Lange:2018pyp; Wofford:2022ykb; Wagner:2025bih; PhysRevD.96.104041. RIFT utilizes a grid-based likelihood calculation approach that aims to converge to the true value over the course of multiple iterations. This formulation allows for parallelization and is therefore substantially faster than standard MCMC or nested sampling based PE pipelines. This is especially useful when using computationally expensive waveforms such as EOB models. Another distinguishing feature of RIFT is that it marginalizes over extrinsic parameters and computes their posteriors in a separate stage. This accelerates the likelihood calculations for intrinsic parameters most relevant to our degeneracy study.
For each system in Table 1, we inject it into the three-detector network of Hanford, Livingston, and Virgo at their design sensitivity LIGO-T2200043 and assuming a zero-noise realization. The extrinsic parameters are given in Table 2, and the distance is adjusted to ensure that all injections have a signal to noise ratio (SNR) of $\rho=30$.
We use broad and agnostic priors on the intrinsic and extrinsic parameters for our eccentric signal model, as described in Appx. A. Since SEOBNRv5EHM is an aligned-spin model, we adopt the aligned-spin prior on the component of the spin along the orbital angular momentum Lange:2018pyp, which tends to favor small spins. We set the prior on eccentricity $e_{\mathrm{10}}$ and relativistic anomaly $\zeta$ to be uniform with $e_{\mathrm{10}}\in[0,0.35]$ and $\zeta\in[0,2\pi)$ at the reference frequency.
@@ IV.2 Results of injection-recovery
In Fig. 5 we see that for the majority of our injections, the recovery favors zero eccentricity. There’s only a singular case where the marginalized posterior $p(e_{\mathrm{10}}\mid d)$ clearly peaks away from $e_{\mathrm{10}}=0$. Inj #5 in Table 1 is our most significant case of misidentification. It is an equal-mass, precessing system of total mass $250\,M_{\odot}$ with $\chi_{1,x}=\chi_{2,x}=0.7$, resulting in $\chi_{p}=0.7$ and $\chi_{\mathrm{eff}}=0$. We can use these marginalized posteriors to evaluate the Bayes factor $\mathcal{B}_{E/C}$ comparing the hypothesis of nonzero eccentricity to the quasi-circular limit. For this we invert the Savage-Dickey density ratio Dickey:1971; Verdinelli:1995, giving
@@ V Discussion
This work points to multiple directions for better understanding and mitigating the possible misidentification of eccentricity. The sensitivity of this degeneracy to the morphology of the signal, for example through varying inclination and precessional dynamics, means that a wider (and computationally expensive) exploration of parameter space is required to better understand it. Ultimately, self-consistent models which incorporate both eccentricity and spin-precession are required to fully account for such degeneracies and measure the properties of astrophysical BBHs. Complete inspiral-merger-ringdown models with such effects are only just becoming available Gamba:2024cvy. The accurate development of these and future models relies heavily on the availability of NR waveforms that capture these physical effects. Therefore, the expansion of NR catalogs especially with higher eccentricities, higher precession, and well resolved higher modes are needed to increase robustness as well as expand the domain of validity for waveform models.
@@ Appendix A Parameter conventions
We mostly use standard conventions in GW astronomy, e.g. LIGOScientific:2025hdt. Two important exceptions are that we adopt a mass ratio $q\geq 1$ as commonly used in the NR community, and we report detector frame masses in all cases, without undoing the effects of cosmic redshifting of the signals. This is appropriate given that our focus is on signals with a small number of cycles in the detector’s sensitive frequency band, and it is the detector frame (rather than true source-frame) parameters that determine the signal as observed.
@@
The dimensionless spins are $\vec{\chi}_{i}$, and we decompose them in a Cartesian coordinate system where the binary angular momentum is along the $z$ axis and the $x$ and $y$ axes span the orbital plane, with the components lying on the $x$-axis at the reference frequency. Conventionally two combinations of the spins are reported due to their importance in the overall evolution of the signal. The effective aligned spin Ajith:2009bn; Santamaria:2010yb is defined as
@@
where $\vec{\chi}_{i}$ are the dimensionless spin vectors of the two components and $\hat{L}$ is the unit vector along the orbital angular momentum. The effective spin is approximately conserved during inspiral Gerosa:2015tea, and encodes the leading spin-effects on the GW signal, at least at comparable masses.
The effective precession parameter $\chi_{p}$ quantifies the in-plane spin components that drive precession Schmidt:2014iyl and is defined as

### v2

@@ I Introduction
This past year marked the tenth anniversary of the first detection of gravitational waves (GW) from the merger of a binary black hole (BBH). The loud chirp of GW150914 Abbott and others (2016) was a landmark discovery that propelled the community into various streams of scientific investigation. Since that first discovery, the The LIGO-Virgo-Kagra (LVK) collaboration has reported 218 events in the four gravitational-wave catalogs Abbott and others (2019, 2021, 2024, 2023); Abac and others (2025e). As the number of observed events increases, the likelihood of encountering systems with non-generic or exotic features, such as orbital eccentricity and spin-induced precession also increases Hoy et al. (2025); Gumbel (1958).
These non-generic features, in turn, offer insight into one of the key open questions in the field: the identification of BBH formation mechanisms; How and when the binary formed, and in what environment. Binary properties like eccentricity or spins tilted below the orbital plane can indicate that the binary underwent dynamical interactions at some point in its lifetime, e.g. Samsing (2018); Rodriguez et al. (2016). Thus, accurate and reliable measurement of these features is of great importance. Occasionally, the GWs observed by the detectors will capture only the final cycles of a binary system’s inspiral and merger. That can give rise to the challenge of using limited information, sometimes only a handful of cycles in the sensitive band, to reconstruct as much of the binary’s history as possible.
The GW observations have revealed a potential population of massive BHs Abac and others (2025d), with examples like GW190521 having an estimated total mass of $\sim 150\,M_{\odot}$ Abbott and others (2020a, b) and GW231123 with $\sim 240\,M_{\odot}$ Abac and others (2025e, a). These massive signals have few GW cycles in band, only $\sim 4$–5, making robust interpretation challenging Fumagalli et al. (2024). A small number of cycles poses a challenge in placing constraints on the source properties, particularly parameters such as eccentricity and in-plane spin components. The initial LVK analysis of GW190521 included quasi-circular precessing templates Abbott and others (2020a, b), and found mild evidence for the imprint of precession; however, it did not include any with eccentricity. Since then, multiple studies have reanalyzed the event; however, its origin remains ambiguous. Work by various groups has shown that the signal could have been emitted by an eccentric source Romero-Shaw et al. (2020); Gayathri et al. (2022), a precessing source Miller et al. (2024), or more exotic compact objects such as boson stars Calderón Bustillo et al. (2021) or cosmic strings Cuceu et al. (2025). These competing propositions highlight the challenge of positively inferring eccentricity from short signals, and the possibility that precession and eccentricity can be misidentifed with limited models.
@@
Eccentricity and spin-precession are relatively well-studied parameters individually. Both these effects cause a modulation in amplitude and frequency in the GW signal. However, these effects generally operate on different timescales. While eccentricity produces modulations at an orbital timescale, precession can cause modulations on longer timescales, spanning many orbits. Superficially, these distinct timescales can make the waveforms produced by systems displaying eccentric or precessing dynamics appear starkly different. Nevertheless, as illustrated in Fig. 1, there exists regions of parameter space in which their modulations overlap in amplitude and phase in a way that can make them qualitatively similar and potentially degenerate. Romero-Shaw et al. Romero-Shaw et al. (2023) reported that a sufficient number of cycles in-band provides enough information to distinguish between these effects; however, this is not the case for a GW190521-like signal, i.e., signals with few cycles in-band. Contrary to expectations, Xu et al. Xu and Hamilton (2023) focused on heavier systems but reported no correlation between eccentricity and precession in the specific numerical relativity (NR) systems they studied. More recently, Divyajyoti et al. Divyajyoti and others (2025) expanded this picture by exploring parameter biases for eccentric systems with and without precessional effects when analyzed with waveform models with incomplete physics. They find that for systems exhibiting both these effects simultaneously, an eccentric aligned spin model is preferred over a quasi-circular precessing model. This highlights how waveform systematics can mimic an underlying physical degeneracy. Further investigating these effects, Gupta et al. Gupta et al. (2025) show posteriors from two distinct binary systems where eccentricity is misidentified as spin-precession. Their results point to the existence of the degeneracy uni-directionally. However, the overall prevalence, parameter dependence, and strength of this degeneracy in the broader parameter space remain largely unclear.
@@
Meanwhile, the astrophysical importance of measuring eccentricity has driven developments in the modeling of GWs from eccentric systems, as well as assessing the impact of mismodeling on PE. A number of state-of-the-art models exist that treat precession and eccentricity separately (see Sec. II). Such models show that neglecting eccentricity when present leads to biases in interpretations, e.g. O’Shea and Kumar (2023); Favata et al. (2022); Divyajyoti et al. (2024); Saini et al. (2022); Zeeshan and O’Shaughnessy (2024). They have been used to supplement quasi-circular, precessing inferences of detected BBH GW events with additional eccentric analyses, e.g. Romero-Shaw et al. (2021); Iglesias and others (2024); Gupte and others (2024); Planas et al. (2025); Romero-Shaw et al. (2025), which in some cases suggest non-negligible eccentricities for a subset of events. Finally, a small number of waveform models have been developed which incorporate both eccentricity and precession Klein et al. (2018); Gamba et al. (2024a); Morras et al. (2025b, a). These are either limited to only the inspiral phase, or are relatively new models which remain computationally expensive for use in full-scale PE studies.
@@
This paper is organized as follows. We describe the waveform model and associated conventions in Sec. II. In Sec. III, we present the setup and results of the preliminary mismatch study that informs our parameter estimation section. In Sec. IV, we detail the setup and results for PE for the identified injection set. Finally, in Sec. V, we discuss the implications and robustness of our results and outline directions for future work. Throughout this work we set $G=c=1$, quote detector frame rather than (redshifted) source-frame masses Abac and others (2025b), and adopt the NR convention that the mass ratio is greater than unity, $q\geq 1$.
@@ II Waveform model
A number of state-of-the-art waveform models exist with differing conventions, approximations, and physics content. This includes phenomenological models Pratten et al. (2021); Colleoni et al. (2025); Hamilton et al. (2021); Thompson et al. (2024); Hamilton et al. (2025); Estellés et al. (2022b, a), effective-one-body (EOB) models from the SEOB family Bohé et al. (2017); Cotesta et al. (2018); Ossokine et al. (2020); Cotesta et al. (2020); Mihaylov et al. (2021); Pompili et al. (2023) and the TEOB family Nagar et al. (2020); Riemenschneider et al. (2021); Akcay et al. (2021); Gamba et al. (2022, 2024b), and the NR surrogate models Varma et al. (2019a, b). Many of these model families also have there own associated eccentric model de Lluc Planas et al. (2025); Gamboa et al. (2024); Ramos-Buades et al. (2022); Chiaramello and Nagar (2020); Islam et al. (2021). For this work, we use SEOBNRv5 waveforms that are based on the EOB formalism. EOB models approximate the full inspiral–merger–ringdown regime. We adopt SEOBNRv5EHM Gamboa et al. (2024) as our non-precessing eccentric model and SEOBNRv5PHM Ramos-Buades et al. (2023); Pompili et al. (2023) as our quasi-circular precessing model. Using waveforms from the same model family minimizes potential model-related systematics when comparing the two. Both models support higher-order modes. The modes used for SEOBNRv5EHM are {(2,±2), (2,±1), (3,±3), (3,±2), (4,±4), (4,±3)}, and for SEOBNRv5PHM are {(2,±2), (2,±1), (3,±3), (3,±2), (4,±4), (4,±3), (5,±5)} in the co-precessing frame. SEOBNRv5EHM allows for explicit sampling in relativistic anomaly $\zeta$ at the reference frequency, an essential parameter in the relativistic definition of eccentricity.
@@ III.1 Parameter choices
Mismatches Owen and Sathyaprakash (1999) are conventionally used to quantify the similarity between two waveforms, for example when calibrating analytical models against post-Newtonian or NR waveforms Boyle et al. (2007); Hannam et al. (2008); Buonanno et al. (2009); Ajith and others (2007). In this work, we use the mismatch between waveforms generated by physically distinct systems to assess the similarity between eccentric and precessing signals. This serves as a relatively inexpensive diagnostic for the expected behavior of full PE. This is a suboptimal approach, as we can only maximize over a few parameters while keeping several extrinsic and intrinsic parameters fixed.
@@
where $h_{p}$ is the precessing signal, $h_{e}$ is the eccentric signal and $\langle\cdot|\cdot\rangle$ is the standard noise-weighted inner product (Appx. B). For the mismatch we maximize the inner product over the difference in coalescence time $\Delta t$, coalescence phase $\Delta\phi$, and polarization angle $\Delta\psi$ between the signals. We fix the same total mass $M_{\mathrm{tot}}$ and mass ratio $q$ for both signals. In all cases we fix the inclination to $\pi/3$ radians, a value intermediate between the special face-on ($\iota=0$) and edge-on ($\iota=\pi/2$) cases. This inclination value also coincides with the peak of the expected observed distribution of inclinations Schutz (2011). Meanwhile, as the two models support different physical effects, our choices of spins and eccentricity parameters necessarily differ.
@@
The number of cycles in band is a significant factor in identifying eccentricity Romero-Shaw et al. (2023). For a given total mass, the mass ratio and number of cycles are inversely related Cutler and Flanagan (1994) (see Fig. 2); therefore we expect that misidentification of eccentricity is more likely in short duration signals. As a result, we focus on high total mass systems having $M_{\text{tot}}\in[200M_{\odot},300M_{\odot}]$ (detector-frame), while varying the mass ratios $q\in[1,3]$ to control the number of cycles in band.
@@ IV.1 PE methodology
We use Bayesian inference on injected signals using the parameter inference pipeline RIFT Pankow et al. (2015); Lange et al. (2018); Wofford and others (2022); Wagner et al. (2025); Lange et al. (2017). RIFT utilizes a grid-based likelihood calculation approach that aims to converge to the true value over the course of multiple iterations. This formulation allows for parallelization and is therefore substantially faster than standard MCMC or nested sampling based PE pipelines. This is especially useful when using computationally expensive waveforms such as EOB models. Another distinguishing feature of RIFT is that it marginalizes over extrinsic parameters and computes their posteriors in a separate stage. This accelerates the likelihood calculations for intrinsic parameters most relevant to our degeneracy study.
For each system in Table 1, we inject it into the three-detector network of Hanford, Livingston, and Virgo at their design sensitivity 57 and assuming a zero-noise realization. The extrinsic parameters are given in Table 2, and the distance is adjusted to ensure that all injections have a signal to noise ratio (SNR) of $\rho=30$.
We use broad and agnostic priors on the intrinsic and extrinsic parameters for our eccentric signal model, as described in Appx. A. Since SEOBNRv5EHM is an aligned-spin model, we adopt the aligned-spin prior on the component of the spin along the orbital angular momentum Lange et al. (2018), which tends to favor small spins. We set the prior on eccentricity $e_{\mathrm{10}}$ and relativistic anomaly $\zeta$ to be uniform with $e_{\mathrm{10}}\in[0,0.35]$ and $\zeta\in[0,2\pi)$ at the reference frequency.
@@ IV.2 Results of injection-recovery
In Fig. 5 we see that for the majority of our injections, the recovery favors zero eccentricity. There’s only a singular case where the marginalized posterior $p(e_{\mathrm{10}}\mid d)$ clearly peaks away from $e_{\mathrm{10}}=0$. Inj #5 in Table 1 is our most significant case of misidentification. It is an equal-mass, precessing system of total mass $250\,M_{\odot}$ with $\chi_{1,x}=\chi_{2,x}=0.7$, resulting in $\chi_{p}=0.7$ and $\chi_{\mathrm{eff}}=0$. We can use these marginalized posteriors to evaluate the Bayes factor $\mathcal{B}_{E/C}$ comparing the hypothesis of nonzero eccentricity to the quasi-circular limit. For this we invert the Savage-Dickey density ratio Dickey (1971); Verdinelli and Wasserman (1995), giving
@@ V Discussion
This work points to multiple directions for better understanding and mitigating the possible misidentification of eccentricity. The sensitivity of this degeneracy to the morphology of the signal, for example through varying inclination and precessional dynamics, means that a wider (and computationally expensive) exploration of parameter space is required to better understand it. Ultimately, self-consistent models which incorporate both eccentricity and spin-precession are required to fully account for such degeneracies and measure the properties of astrophysical BBHs. Complete inspiral-merger-ringdown models with such effects are only just becoming available Gamba et al. (2024a). The accurate development of these and future models relies heavily on the availability of NR waveforms that capture these physical effects. Therefore, the expansion of NR catalogs especially with higher eccentricities, higher precession, and well resolved higher modes are needed to increase robustness as well as expand the domain of validity for waveform models.
@@ Appendix A Parameter conventions
We mostly use standard conventions in GW astronomy, e.g. Abac and others (2025b). Two important exceptions are that we adopt a mass ratio $q\geq 1$ as commonly used in the NR community, and we report detector frame masses in all cases, without undoing the effects of cosmic redshifting of the signals. This is appropriate given that our focus is on signals with a small number of cycles in the detector’s sensitive frequency band, and it is the detector frame (rather than true source-frame) parameters that determine the signal as observed.
@@
The dimensionless spins are $\vec{\chi}_{i}$, and we decompose them in a Cartesian coordinate system where the binary angular momentum is along the $z$ axis and the $x$ and $y$ axes span the orbital plane, with the components lying on the $x$-axis at the reference frequency. Conventionally two combinations of the spins are reported due to their importance in the overall evolution of the signal. The effective aligned spin Ajith and others (2011); Santamaria and others (2010) is defined as
@@
where $\vec{\chi}_{i}$ are the dimensionless spin vectors of the two components and $\hat{L}$ is the unit vector along the orbital angular momentum. The effective spin is approximately conserved during inspiral Gerosa et al. (2015), and encodes the leading spin-effects on the GW signal, at least at comparable masses.
The effective precession parameter $\chi_{p}$ quantifies the in-plane spin components that drive precession Schmidt et al. (2015) and is defined as